Assume that you are in the 25 percent marginal tax bracket and that you have $5,000 to invest. You have narrowed your investment choices down to California bonds with a yield of 7 percent or equally risky ExxonMobil bonds with a yield of 10 percent. Which one should you choose and why? At what marginal tax rate would you be indifferent to the choice between California and ExxonMobil bonds?
Answer #1
After tax rate = YTM * (1-Tax rate)
After tax rate = 10 * (1-0.25)
After tax rate = 7.5

Choose Exxon bond as it has higher after tax yield

After tax rate = YTM * (1-Tax rate)
7 = 10 * (1-Tax rate)
Tax rate% = 30 = indifference tax rate